OpenStudy (anonymous):

Help me please . :( solve for x . 7^x-2=19 2lnx-2ln3=0 8-3e^4x=5 log2(x-2)+log2(3=4 4log3(x+2)=1 log4x-log4(x-1)=1/2

OpenStudy (anonymous):

\[7^x-2=19 \rightarrow 7^x=19+2 \rightarrow 7^x=21\] Taking log both sides, we find \[\log7^x= \log 21\rightarrow x \log7= \log (7 \times 3)\] \[\rightarrow x \log7= \log 7 + \log 3 \rightarrow x \times 0.8451= 0.8451 + 0.4771\] \[ \rightarrow x \times 0.8451= 1.3222 \rightarrow x = \frac{1.3222}{0.8451} =1.5645\] Hence x=1.5645 @JustTooMuchh

OpenStudy (anonymous):

2lnx-2ln3=0 i.e. 2ln x-2ln 3=ln 1 [since In 1 = 0] \[\ln x^2-\ln 3^2=\ln 1 \rightarrow \ln x^2-\ln 9=\ln 1 \rightarrow \ln \frac {x^2}{9}=In 1\] \[\frac {x^2}{9}=1 \rightarrow x^2 = 9 \rightarrow x = \pm 3\] but \[-3 \in x\] therefore x =3 @JustTooMuchh